OpenStudy (anonymous):

The number of cars (c) in a parking lot increases when the parking fee (f) decreases. Write the correct equation for this scenario, and solve for the number of cars when the fee is $6. Fee Cars 20 15 30 10 c = − 300 over f c = 300 over f c = 3f; c = 18 c = 3f; c = 2 Can anyone please guide me through this question? @johnweldon1993 @ajprincess @phi @Opcode @Gabylovesyou @anonymous_user @DLS

OpenStudy (anonymous):

@myininaya @Compassionate

OpenStudy (anonymous):

@zepdrix can u plz help out?

OpenStudy (anonymous):

@KlOwNlOvE @MrNood @midhun.madhu1987

OpenStudy (anonymous):

can anyone just help me understand what I need to do? i need help!

OpenStudy (johnweldon1993):

Well lets see...when 1 function increases, the other decreases...that reminds me of an inverse variation...or \[\large c = \frac{k}{f}\] As that function shows, as F gets bigger, C gets smaller...but what we want (and have) is that as F (the fee) gets smaller) the cars (C) gets bigger. So now...we need to be able to solve for that 'k' there....so lets choose a point we have from that little table there

OpenStudy (johnweldon1993):

when there are 10 cars...the fee was 30 dollars...so \[\large c = \frac{k}{f}\] \[\large 10 = \frac{k}{30}\] \[\large k = 300\] so there will be a constant of variation of 300....so our new quation that we will want to use to solve our problem will be \[\large c = \frac{300}{f}\]

OpenStudy (johnweldon1993):

But just to be sure...lets check the second point we have on the table 15 cars at a fee of 20 dollars. so \[\large 15 = \frac{300}{20}\] \[\large 300 = 300\] so yes it checks out...

OpenStudy (johnweldon1993):

Now, we are to solve how many cars there will be when the fee (f) is 6 dollars...so we plug in 6 for 'f' in our equation and solve for 'c' \[\large c = \frac{300}{6}\] \[\large c = ?\]

OpenStudy (anonymous):

c= 50

OpenStudy (johnweldon1993):

Perfect...there will be 50 cars when the fee is just 6 dollars

OpenStudy (anonymous):

OOOOOHHH! okaay but why did we choose the second one we couldve plugged in 6 for f on the other choices and get a random answer @johnweldon1993

OpenStudy (johnweldon1993):

Not sure what you mean...

OpenStudy (anonymous):

can u explain to me why it had to be choice B out of all the choices? sorry i'm confused myself lol

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

But we just solved for why it would be choice B Note that with the little table we were given, we can just take the values given and plug them in to check if we didnt want to do all the work so for example with the 15 cars at a fee of 20 dollars... if we plug that into the first choice we would have \[\large 15 = \frac{-300}{20}\] \[\large 300 \cancel{=}-300\] so that choice wouldnt have worked and do it again with the other choices etc... easy way to guess and check instead of doing the work above

OpenStudy (anonymous):

OHHHHH okayyy noooww i get itt!! thanks soo much!

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

why did u use 15 and 20 when they say that fees is 6?

OpenStudy (anonymous):

i'm slow in math LOL

OpenStudy (anonymous):

So can is B the answer?

OpenStudy (anonymous):

@johnweldon1993